OpenStudy (gracygirl):

Fan & Medal Write the point-slope form of the given line that passes through the points (0, -3) and (4, 1). Identify (x1, y1) as (0, -3). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

OpenStudy (error1603):

welcome

OpenStudy (gracygirl):

@Will.H

OpenStudy (will.h):

we have the following points (0,-3) (4,1) use slope formula 1st to find the slope y2-y1/x2-x1 1 + 3/4-0 4/4 = 1 so the slope is equal to 1 now we will use y - y1 = m(x - x1) they told us to pick the point (0,-3) as y1 and x1 so substitute in the original expression and note that m is the slope y - (-3) = 1(x - 0) y + 3 = x - 0 that's the point slope form hope it helps